L=2W
A=50ft^2
what is the perimeter?

Found 3 solutions by Tobiasz, Alan3354, richwmiller:
Answer by Tobiasz(54)

2w + 50ft = P
P = 15

Answer by Alan3354(69443)

Area = L*W
Area = W*2W = 50
w^2 = 25
W = 5
L = 10
Perimeter = 2L+2W = 30

Answer by richwmiller(17219)

L*W=50
L=2W
2w^2=50
w^2=25
w=5
L=10
10+5+10+5=30=Perimeter